OpenStudy (anonymous):

Given the equation 4x + 2y = 9, which equation below would cause an inconsistent-independent system? 6x + 12y = 13 2x + y = -6 x - 2y = 7 8x - 4y = -5

OpenStudy (anonymous):

inconsistent means there's no shared point on the lines aka the lines are parallel

OpenStudy (anonymous):

You mean there isn't going to be any solution?

OpenStudy (anonymous):

the lines in an inconsistent system will not get any solution we pick the line that has no solution with line in problem statement

OpenStudy (anonymous):

I don't know which one that is.

OpenStudy (anonymous):

8x - 4y = -5 <- this one isn't it, the slope is different 2x + y = -6 (*2) 4x + 2y = -12 ^ this one looks exactly the same like the one in the problem statement it has a different solution even though it has the same amounts of y, x that means there's a contradiction and the system is inconsistent

OpenStudy (anonymous):

Where did you get 4x + 2y = -12 from?

OpenStudy (anonymous):

Is it an example?

OpenStudy (anonymous):

from 2x + y = -6 I multiplied it by 2

OpenStudy (anonymous):

LHS and RHS - every term x2

OpenStudy (anonymous):

Ohhh

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

So that's the answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

no problem.

OpenStudy (anonymous):

whats the answer

OpenStudy (anonymous):

@realmadridCR7 inconsistent means that there is never a shared point. the x-y pair the one needs doesn't work in the other. there is not a single x-y pair that both equations will satisfy. so the system of equations can never be satisfied, the system is inconsistent. 

OpenStudy (anonymous):

dependent means ,that the second equation is just a "derived form" of the first, so that second equation "depends" on the first one, example: y= 2x + 3 #first equation 2y = 4x + 6 #second equation The second line is the first line multiplied by two. It is not a new arrangement - it is just multiplied by 2, and it is the first one multiplied by two, so it DEPENDS on the first one.

OpenStudy (anonymous):

you have to compare this equation 4x + 2y = 9 with all the other equations, and determine if there is one that will be parallel; and if it is parallel, then the second requirement is that it's not a multiple of 4x + 2y = 9.